# Disclosure Avoidance for the 2020 Census Demographic and Housing Characteristics File

### v1

@@ 2.2. DHCP Schema
For the persons data universe, the DHCP schema is defined by the following attributes and levels.
@@
  1. (Inv1)
    The DHCP MDF is constrained to be consistent with the persons redistricting data, e.g., any tabulation in the redistricting data file will be the same as this tabulation derived using the DHCP MDF.
  2. (Inv2)
@@
Relationship to Householder or GQ Type
Age Range
Householder
15-115
Spouse or unmarried partner of householder
15-115
Adopted or Biological child
0-89
Stepson/stepdaughter
0-89
Brother or sister
0-115
Parent
30-115
Grandchild
0-74
Parent-in-law
30-115
Son-in-law/daughter-in-law
15-89
Other relative
0-115
Housemate/roommate
0-115
Foster child
0-20
Other non-relative
0-115
Non-military correctional facilities (GQ Types 101-105)
15-115
Group home/correctional/treatment facilities for juveniles (GQ Types 201-203)
0-25
Nursing facilities (GQ Type 301)
20-115
Hospitals/hospices (GQ Types 401-403)
0-115
Residential schools for people with disabilities (GQ Type 405)
3-30
College/university student housing (GQ Type 501)
16-65
Military Housing (GQ Types 106, 404, 601, 602)
17-65
Adult group homes/treatment centers (GQ Types 801, 802)
16-115
Maritime/merchant vessels (GQ Types 900)
16-75
Other non-institutional (GQ Types 701-706, 901-904)
0-115
@@ 2.3. Types of Constraints Within the DAS
For both the DHCH and DHCP DAS implementations, it is worthwhile to separate constraints into two categories. First, edit and structural zero constraints are imposed on the output of the DAS to ensure that the constraints imposed during the creation of the CEF are also satisfied in the output MDF. These constraints can be viewed as the requirements provided by subject matter experts for the face validity of the microdata. These constraints are also data-independent and publicly available. As a result, and also since our privacy guarantees are specified with respect to the records in the CEF, these constraints do not limit the types of inferences on the CEF records that we protect against. Some examples of constraints in this category include that each householder record in the DHCP MDF is constrained to be at least 15 years of age, and, in the DHCH MDF, that each household of size one with a householder with age less than 60 is constrained to have ELDERLY indicate that no household member is over 60 years of age. Other examples of these constraints include constraints (C3-C7) and (Inv3) in the list of DHCH and DHCP constraints above, respectively.
Second, our production DHC implementations also impose several constraints, which we call invariant constraints, that are dependent on the input data files of the DAS DHC implementations rather than on publicly-available data. Specifically, the invariant constraints imposed on the production DHC implementations are (A) the total populations of each state, Washington D.C., and Puerto Rico, (B) the tabulation block locations of each occupied GQ by detailed GQ type, e.g., the location of each nursing facility, and (C) the tabulation block locations of each housing unit. The first set of constraints is imposed in the persons universe of the production 2020 Redistricting Data File DAS implementations. This is carried forward into the production DHCP implementation through our constraints that ensure the DHC outputs are consistent with the redistricting data file. While we protect inferences on whether a given a housing unit is occupied or vacant, we do not provide this same protection to GQs because unoccupied GQs are out of scope for the 2020 Census and the DAS has no access to that information. More detail on the rationale for the inclusion of these invariant constraints is provided by [14].
@@ 2.3.1. Unused Constraints in DHC
Note that the constraints imposed within the DHC DAS implementation do not avoid all inconsistent counts. For example, since the persons and units universe are processed separately from one another, there are some blocks with occupied housing units in the DHCH MDF but without any population in the DHCP MDF. Inconsistent counts such as these are often a feature of sparsely populated areas, as these areas often pose a higher risk of disclosure. The presence of inconsistent counts is a warning to the user about the presence of disclosure avoidance. It is worth noting that inconsistent counts are an accepted feature in other countries, such as in the UK Census [21, 24].
@@ 2.3.2. How Constraints are Imposed
As described previously, the primary way that fitted histograms are computed to satisfy the DHCH and DHCP constraints is through a series of least squares and mixed integer programming optimization solves. These optimization problems are described in detail by [14]. As is also described in [14], we require the constraint matrix for the MIP optimization problems to satisfy a condition known as total unimodularity. This restriction ensures these MIP solves are computationally tractable; for more detail on total unimodularity and the sufficient condition we use to ensure it, see [14, 15].
@@
With the exception of the DHCP constraints described in (Inv2) from Section 2.2, i.e., the DHCP upper and lower GQ count constraints, all other DHCH and DHCP constraints are imposed using these optimization solves. A different approach is required for DHCP constraints described in (Inv2) because the sufficient condition we use to ensure total unimodularity would not hold if these constraints were added to these MIP optimization problems. For this reason, we created a post-processing method called the DHCP optimization heuristic that improves the feasibility for the DHCP constraint (Inv2) of the histogram estimate that is output from the least squares and MIP optimization solves. The DHCP optimization heuristic works by iteratively choosing pairs of histogram elements within the same histogram of a single geographic unit and then considering increasing one count by one and decreasing the other by one. If changing the histogram counts in this way improves feasibility of a DHCP constraint in (Inv2) and does not result in any of the remaining constraints being infeasible, this change to these two histogram counts is accepted, and otherwise the change is rejected and no change is made to these two histogram counts. Afterward, the algorithm goes on to consider the next change in a pair of histogram counts.
In more granular geographic units, i.e., optimized block groups and census blocks, the DHCP optimization heuristic chooses pairs of histogram cells for which all attributes other than RELGQ are identical. Since it is possible for one of the DHCP constraints described in Section 2.2 to become infeasible for all such changes to pairs of histogram counts, the DHCP constraint described in (Inv2) does not always hold for each of these geographic units, as described in the next example.
Example 1.
Suppose the census block histogram estimate that is input into the DHCP optimization heuristic encodes one person who is 31 years old that lives in a hospital, hospice, or treatment facility (GQ types 401-404), but no people residing in residential schools for people with disabilities (GQ type 405). Also suppose the DHCP upper and lower GQ count constraints described in (Inv2) require that at least one respondent resides in GQs of type 405 and that no respondents reside in GQs of types 401-404. In order to maintain feasibility for the DHCP constraints described in (Inv1), i.e., the constraints that ensure consistency with the redistricting data file MDF, we can only consider switching the respondent who currently has a GQ type of 401-404 to a GQ type of 405, since the number of respondents estimated to be in GQ types 401-405 is held fixed by this DHCP constraint. However, switching the GQ type of this respondent from a GQ type in the range 401-404 to 405 would impair the feasibility of the DHCP constraints descried in (Inv3) because this constraint requires that respondents with a GQ type of 405 be less than 30 years old. Since all changes that we consider for this block geographic unit result in the feasibility of either the DHCP constraint described in (Inv1) or in (Inv3) to fail, the DHCP optimization heuristic would not make any changes to the input histogram estimate in this case.
However, in the case of all geographic units that are less granular than optimized block groups, we also consider switching respondents between pairs of histogram cells with different values for the attribute AGE. As described in the next example, this results in all of the DHCP constraints described in (Inv1-Inv3) being satisfied for each geounit.
Example 2.
We consider the same setting as Example 1, but, rather than considering the case in which the geographic unit is a census block, we suppose it is a state geounit instead. In this case the DHCP optimization heuristic considers changes to both the RELGQ attribute and the AGE attribute. As described in Example 1, only changes to RELGQ that result in the respondent’s GQ type being changed to GQ type 405 are considered. Since one attribute in the redistricting data file is VOTING_AGE, which is one if the respondent is at least 18 years old and zero otherwise, the DHCP constraints described in (Inv1) require that only changes to the AGE attribute that change the age of the respondent from 31 to any age of at least 18 are considered. In the case of this state geographic unit, the DHCP optimization heuristic considers pairs of histogram counts until it eventually accepts a change that results in this record’s GQ type being switched to 405 and its age being switched to a value between 18 and 30.
@@ 3. Computational Requirements for the DHC DAS
[14] provide detail on the computational environment used by the redistricting data file DAS implementation, including a list of software dependencies, AWS cluster configurations, and Spark settings. In this section, we describe the computational environment used for our DHC DAS implementations, with a particular focus on the aspects that we changed after publication of [14].
@@ 4. DHC Accuracy Experiments and PLB Allocations
In preparation for the 2020 DHC production DAS implementations, the DAS team performed a series of accuracy experiments using 2010 Census inputs in order to support recommendations for the Census Bureau’s final production settings of the 2020 DHC production implementations, including the PLB allocations. The DHC accuracy experiments were similar to those used to support the final production settings for the 2020 Redistricting Data (P.L. 94-171) Summary File production DAS implementations [14]. Specifically, these accuracy experiments used an iterative process consisting of DAS configuration updates based on input from subject-matter experts (SMEs) on the accuracy of tabulations from prior experiment iterations. The Census Bureau internal SMEs included those from the Population Division, and Social, Economic, and Housing Statistics Division (SEHSD).
@@
External SMEs also provided feedback based on the output of two DHC demonstration data products (DDPs) based on 2010 Census input data released on March 16, 2022 and August 25, 2022, respectively. [^5] Despite the name “demonstration data products,” these releases were not meant to demonstrate the final production settings we intended to use. These were instead intended to demonstrate performance of the DHC DAS implementation for the settings we were currently testing at the time in our DHC accuracy experiments. This includes feedback provided through a National Academies of Science, Engineering, and Medicine’s Committee on National Statistics (CNSTAT) meeting in June of 2022 [18], academic conferences, and external stakeholders independently contacting the Census Bureau regarding the accuracy of these 2010 DDPs for specific use cases. In addition to these two DDPs, the DAS output based on 2010 Census input data and using the 2020 DHC production implementation settings was also released on April 3, 2023. All of these releases are currently available at [26].
@@
The result of these accuracy experiments was the 2020 DHC production configuration parameters, including strategy matrices for both DHCH and DHCP, which refer to the marginals that are chosen to become noisy measurements as well as the PLB allocation of each marginal that defines the scale of the noise added to the marginal for each geographic level. As described in [14] in more detail, we use the $\rho-$zero-concentrated differential privacy ($\rho-$zCDP) accounting framework to specify these PLB values [16]. Each strategy matrix described in the next two subsections provides the initial PLB allocations for each geographic unit on the spine, prior to the spine optimization routines’ alteration of these allocations. Specifically, as described by [19] in more detail, in the case of parent and child geographic units for which the parent geographic unit has only one child, the spine optimization routines alter the initial PLB allocations of these geographic units. Aside from dictating the scale of the noise for each noisy measurement used within the DAS, the PLB allocations are also useful because they can be used to derive upper bounds on the power of certain hypothesis tests on any given individual record in the CEF; for more detail, see [23].

### v2

@@ 3.2. DHCP Schema
Unlike the DHCH DAS executions, there is only one category of counts that are estimated in DHCP DAS executions, which are counts of people, including both people in housing units and GQs, so DHCP DAS executions only use one histogram. Specifically, the DHCP schema is defined by the following attributes and levels.
@@
  1. (P1)
    The DHCP MDF is constrained to be consistent with the persons redistricting data, e.g., recomputing any tabulation that is in the redistricting data file using the DHCP MDF results in the same values as those in the redistricting data file.
  2. (P2)
@@
Relationship to Householder or GQ Type Age Range
Householder 15-115
Spouse or unmarried partner of householder 15-115
Adopted or Biological child 0-89
Stepson/stepdaughter 0-89
Brother or sister 0-115
Parent 30-115
Grandchild 0-74
Parent-in-law 30-115
Son-in-law/daughter-in-law 15-89
Other relative 0-115
Housemate/roommate 0-115
Foster child 0-20
Other non-relative 0-115
Non-military correctional facilities (GQ Types 101-105) 15-115
Group home/correctional/treatment facilities for juveniles (GQ Types 201-203) 0-25
Nursing facilities (GQ Type 301) 20-115
Hospitals/hospices (GQ Types 401-403) 0-115
Residential schools for people with disabilities (GQ Type 405) 3-30
College/university student housing (GQ Type 501) 16-65
Military Housing (GQ Types 106, 404, 601, 602) 17-65
Adult group homes/treatment centers (GQ Types 801, 802) 16-115
Maritime/merchant vessels (GQ Types 900) 16-75
Other non-institutional (GQ Types 701-706, 901-904) 0-115
@@ 3.3. Types of Constraints Within the DAS
For both the DHCH and DHCP DAS implementations, it is worthwhile to separate constraints into two categories. First, edit and structural zero constraints are imposed on the output of the DAS to ensure that the constraints imposed during the creation of the CEF are also satisfied in the output MDF. These constraints can be viewed as the requirements provided by SMEs for the face validity of the microdata. These constraints are also data-independent and publicly available. As a result, and also since our privacy guarantees are specified with respect to the records in the CEF, these constraints do not limit the types of inferences on the CEF records that we protect against. Some examples of constraints in this category include that each householder record in the DHCP MDF is constrained to be at least 15 years of age, and, in the DHCH MDF, that each household of size one with a householder with age less than 60 is constrained to have ELDERLY indicate that no household member is over 60 years of age. Other examples of these constraints include constraints (C3-C7) and (P3) in the list of DHCH and DHCP constraints above, respectively.
Second, our production DHC implementations also impose several constraints, which we call invariant constraints, that are dependent on the input data files of the DAS DHC implementations rather than on publicly-available data. Specifically, the invariant constraints imposed on the production DHC implementations are (A) the total populations of each state, Washington D.C., and Puerto Rico, (B) the tabulation block locations of each occupied GQ by detailed GQ type, e.g., the location of each nursing facility, and (C) the tabulation block locations of each housing unit. The first set of constraints is imposed in the persons universe of the production 2020 Redistricting Data File DAS implementations. This is carried forward into the production DHCP implementation through our constraints that ensure the DHC outputs are consistent with the redistricting data file. While we protect inferences on whether a given a housing unit is occupied or vacant, we do not provide this same protection to GQs because unoccupied GQs are out of scope for the 2020 Census and the DAS has no access to that information. More detail on the rationale for the inclusion of these invariant constraints is provided by [abowd20222020].
@@ 3.3.1. Unused Constraints in DHC
Note that the constraints imposed within the DHC DAS implementation do not avoid all inconsistent counts. For example, since the persons and units universe are processed separately from one another, there are some blocks with occupied housing units in the DHCH MDF but without any population in the DHCP MDF. Inconsistent counts such as these are most common in sparsely populated areas. Ensuring consistency between the persons and units universes using an approach that is similar to the optimization passes that are currently used in the DAS implementation might not be computationally feasible because this would significantly increase the size of the optimization problems solved within DAS. [^5] The next section also describes the sufficient condition that we use to ensure our MIP solves are computationally tractable. This sufficient condition would also fail to hold if we added the constraints required for consistency between the counts in these two universes to combined optimization solves that estimate and fix the DHCH and DHCP histograms simultaneously. Also, it is worth noting that inconsistent counts are an accepted attribute of census tabulations in other countries, such as in the UK Census [doveApplying, ONS:2023].
@@ 3.3.2. How Constraints are Imposed
As described previously, the primary way that fitted histograms are computed to satisfy the DHCH and DHCP constraints is through a series of least squares and mixed integer programming optimization solves. These optimization problems are described in detail by [abowd20222020]. As is also described in [abowd20222020], we require the constraint matrix for the MIP optimization problems to satisfy a condition known as total unimodularity. This restriction ensures these MIP solves are computationally tractable; for more detail on total unimodularity and the sufficient condition we use to ensure it, see [abowd20222020, asano2003matrix].
@@
With the exception of the DHCP constraints described in (P2), i.e., the upper and lower bound constraints on populations in each GQ, all other DHCH and DHCP constraints are imposed by including them directly in the optimization solves. A different approach is required for DHCP constraints described in (P2) because the sufficient condition we use to ensure total unimodularity would not hold if these constraints were added directly to the MIP optimization problems. For this reason, we created a post-processing method called the DHCP optimization heuristic that improves the feasibility for the constraints in (P2) of the histogram estimate that is output from the optimization passes described in Section 2.1. More specifically, this method “improves feasibility" of the upper and lower bound constraints in (P2) on the detailed GQ population counts by moving each detailed GQ population count closer to the interval of feasible GQ population counts, but it does not ensure that all of these constraints are actually satisfied in all cases, as we describe below in more detail. This method works by iteratively moving records between pairs of histogram cells for each individual child geographic unit, with these pairs defined so that these moves maintain feasibility of the DHCP constraints described in (P1) and (P3).
Before describing this optimization heuristic, some additional background will be helpful. Note that the redistricting data file persons MDF was constrained to satisfy similar bounds on the population within each major GQ type. [^6] The major GQ type code is given by the first digit of the three detailed GQ type code for each detailed GQ type code less than 700, and is equal to 7 for all remaining detailed GQ types. For example, the major GQ type code for detailed GQ type 301, i.e., nursing facilities, is 3, and the major GQ type code for detailed GQ type 900, i.e., maritime/merchant vessels, is 7. To continue to satisfy the constraints described in (P1), i.e., the constraints that ensure consistency with the redistricting data file persons tabulations, while moving the population with a given detailed GQ type toward a feasible interval, we only need to consider moving records between pairs of histogram cells with the same major GQ type but different detailed GQ types. Throughout the discussion below we will consider an example of a geographic unit with one maritime/merchant vessels (GQ type 900, which is in major GQ type 7) and no other GQs. We will also suppose that the histogram estimate output from the optimization passes places only one estimated resident in a GQ type, that this person is placed in an emergency and transitional shelter (GQ type 701, which is also in major GQ type 7), and that this respondent is estimated to be 10 years old. This is an interesting example for a few reasons. First, in this case, the constraints described in (P2) require that between 1 and 99,999 records in this geographic unit reside in GQ type 900 and that no one is residing of GQ type 701, so the histogram estimate output from the optimization passes does not satisfy two constraints in (P2). Second, simply switching the GQ type of the sole record estimated to be in major GQ type 7 from 701 to 900 would result in one constraint on AGE in (P3) failing to hold, since each record in GQ type 900 is required to be at least 16 years old.
The DHCP optimization heuristic is run separately on the histogram estimate of each individual child geographic unit, and the method does not ensure that parent-child consistency is maintained. In the example described above, the optimization heuristic may move the sole estimated GQ resident from GQ type 701 to GQ type 900 within this geographic unit. However, since the method is run separately on the child geographic units of each parent without requiring that the changes to the histograms of each child geographic unit cancel with one another, it is likely that the method changes the count of records in these two GQs in total over all of the child geographic units, so these two counts are not likely to be consistent with the previous estimate of the parent geographic unit. Since the TopDown Algorithm only uses the previously estimated geographic level when estimating the next most granular geographic level (rather than using histogram estimates from all previously estimated geographic levels), the typically-small number of record changes carried out by the optimization heuristic each time it is called has the potential to cause larger changes in tabulations of marginals in less-granular geographic units, such as the US geographic unit, after it is applied to the histograms of each of the more granular geographic levels. Below we will describe the issues caused by not ensuring parent-child consistency, after describing the method in more detail.
This heuristic works by iteratively considering each detailed GQ type that is included as a level of the attribute RELGQ. When the histogram counts for this level of RELGQ are satisfied, i.e., the estimate of individuals in this detailed GQ type are between the required lower and upper bounds, the child histogram is not altered and the next detailed GQ type is considered. Otherwise, the optimization heuristic defines a set of pairs of histogram cells, with one element of each pair called the donor cell and the other the recipient cell, and with each pair defined so that moving a record from the donor cell to the recipient cell would move the population estimate for the detailed GQ type being considered toward the interval of feasible counts and also such that certain other conditions hold, which we describe below in more detail. If there exists at least one such pair of donor and recipient cells that provide a way to move a record without violating any of the remaining DHCP constraints, a pair is chosen uniformly at random from these possible pairs and then a record from the chosen donor histogram cell is moved to the chosen recipient histogram cell. This process is repeated until the histogram estimate satisfies the detailed GQ bound constraint or there does not exist a possible record to move without violating one or more other DHCP constraints, at which point the next detailed GQ bound constraint is considered.
The additional conditions that are used to define the pairs of donor and recipient histogram cells are set to balance a tradeoff between feasibility and accuracy of the tabulations. For example, we could ensure all of the counts satisfy all of the DHCP constraints described in Section 3.2 by setting this condition so that all donor and recipient cells have identical values for all attributes other than RELGQ and AGE. In particular, this goal would require allowing for changes in respondents’ AGE attribute (in addition to changes in RELGQ) to restore feasibility of the constraints described in (P3), in the event that the set of feasible ages for a record change due to change in its RELGQ attribute. Returning to our running example, as described above, if we were to change the GQ type of the sole record estimated to be in a GQ from 701 to 900, then maintaining feasibility of the constraints described in (P3) would require that we also change the AGE attribute of this record to at least 16. After testing this approach of ensuring feasibility of all DHCP constraints, we found that this resulted in too large of a perturbation in the AGE marginals. Since the optimization heuristic does not ensure parent-child consistency, the aggregate impact of the optimization heuristic on the AGE marginal estimate for the US as a whole can be derived by summing over the changes to the AGE marginals made by the optimization heuristic in all geographic units. For this reason, requiring feasibility of all constraints in each geographic unit was especially problematic for the accuracy of the AGE marginal for the US as a whole. To ameliorate this issue, the DHCP production settings require that all donor and recipient record pairs in the optimized block group and block geographic levels have identical levels for every attribute other than RELGQ. In other words, this optimization heuristic does not consider changes to the AGE marginal in these two most granular geographic levels. In terms of our running example, if this child geographic unit is in either of these two most granular geographic levels, the optimization heuristic would not make any changes to the estimated output from the optimization passes, since doing so would require a change in the AGE attribute of this record.
In all geographic levels higher than the optimized block group and block geographic levels on the spine, we first attempted to define the donor and recipient cell pairs in the same way that is done for these two most granular geographic levels, and only resort to changing the AGE levels of records when this approach is not capable of restoring feasibility of all the constraints described in (P2). So, if the child geographic unit in our running example is in these less-granular geographic levels, the optimization heuristic would return a histogram with the sole record estimated to be in a GQ having its GQ type changed from 701 to 900 and its AGE attribute changed from 10 to either 16 or 17 (both with a probability of $0.5).$ Note that the updated AGE attribute of either 16 or 17 results in both the AGE of the residents of GQ type 900 continuing to be feasible, i.e., feasibility of the constraints in (P3), and, since VOTING_AGE is one attribute included in the persons redistricting data file MDF, the histogram continuing to be consistent with the previously estimated persons redistricting data file MDF, i.e., feasibility of the constraints in (P1). In practice, we found that these settings resulted in both the constraints described in (P2) typically being feasible and the AGE and RELGQ marginals of the less-granular geographic units, such as the US as a whole, being sufficiently accurate.
In terms of runtime, this approach is not a major bottleneck in DAS DHCP executions because the number of records moved between cells is typically not high for our production settings. This is a result of several factors, including the fact that the histogram estimates returned by the optimization passes for most child geographic units are guaranteed to satisfy the constraints in (P2) for a few reasons. Specifically, most geographic units do not have any GQs, especially those at granular geographic levels, and the population in each detailed GQ type for these geographic units for the histogram estimates returned from the optimization passes is constrained to their correct value of zero for these geographic units by the constraints in (P1), since major GQ type was estimated in the persons redistricting data file MDF. Also, even when this is not the case, our 2020 DAS DHCP production settings typically only result in modest errors for each detailed GQ type population, and, since the CEF-based GQ type populations satisfy the constraints in (P2) for each geographic unit, the accuracy of the histogram estimate returned by the optimization passes has the effect of limiting the number of records that are moved by the DHCP optimization heuristic.
@@ 4. Computational Requirements for the DHC DAS
[abowd20222020] provide detail on the computational environment used by the redistricting data file DAS implementation, including a list of software dependencies, AWS cluster configurations, and Spark settings. In this section, we describe the computational environment used for our DHC DAS implementations, with a particular focus on the aspects that we changed after publication of [abowd20222020].
@@ 5. DHC Accuracy Experiments and PLB Allocations
In preparation for the 2020 DHC production DAS implementations, the DAS team performed a series of accuracy experiments using 2010 Census inputs in order to support recommendations for the Census Bureau’s final production settings of the 2020 DHC production implementations, including the PLB allocations. The DHC accuracy experiments were similar to those used to support the final production settings for the 2020 Redistricting Data (P.L. 94-171) Summary File production DAS implementations [abowd20222020]. Specifically, these accuracy experiments used an iterative process consisting of DAS configuration updates based on input from SMEs on the accuracy of tabulations from prior experiment iterations. The Census Bureau internal SMEs included those from the Population Division (POP), and Social, Economic, and Housing Statistics Division (SEHSD).
@@
External SMEs also provided feedback based on the output of two DHC demonstration data products (DDPs) based on 2010 Census input data released on March 16, 2022 and August 25, 2022, respectively. [^7] Despite the name “demonstration data products,” these releases were not meant to demonstrate the final production settings we intended to use. These were instead intended to demonstrate performance of the DHC DAS implementation for the settings we were currently testing at the time in our DHC accuracy experiments. This includes feedback provided through a National Academies of Science, Engineering, and Medicine’s Committee on National Statistics (CNSTAT) meeting in June of 2022 [cnstat], academic conferences, and external stakeholders independently contacting the Census Bureau regarding the accuracy of these 2010 DDPs for specific use cases. In addition to these two DDPs, the DAS output based on 2010 Census input data and using the 2020 DHC production implementation settings was also released on April 3, 2023. All of these releases are currently available at [DHCddps].
As was also described by [abowd20222020] in the context of tuning the DAS for the 2020 production redistricting data file DAS executions, a significant portion of the time that went into tuning DAS during the DHC accuracy experiments focused on collecting accuracy targets for DHC use cases provided by internal and external SMEs and then attempting to satisfy these targets with minimal PLB. In other words, one of the main focuses during these accuracy experiments was solving the dual problem to the more standard problem in the formal privacy literature of maximizing data utility such that the PLB is below a prespecified fixed threshold. This is in part because the majority of the feedback we received over the course of these accuracy experiments was related to concerns with data quality rather than confidentiality. A wide variety of accuracy metrics were used over the course of the DHC accuracy experiments, for a variety of queries and sets of geographic entities, including measures of central tendency of the absolute value of residuals (e.g., mean and median absolute error), quantiles of residuals, and measures of absolute relative residuals. The DAS team also used privacy-semantic statements, primarily those based on bounds on the statistical power of hypothesis tests on the attributes of CEF records, to communicate privacy concerns [kifer2022bayesian]. The DAS team also provided DSEP with results from a reconstruction/reidentification attack using the DAS output from our test runs, which is described in more detail by [abowd20232010]. Ultimately, choosing the production DAS settings that balance the tradeoffs between accuracy, privacy, and the quantity of statistics is a policy decision that is the responsibility of DSEP. DSEP chose the 2020 DHC DAS production settings based on background information, accuracy metrics, and recommendations from internal stakeholders that were based on the results of the DHC accuracy experiments.
@@
The result of these accuracy experiments was the 2020 DHC production configuration parameters, including the DHCH and DHCP strategy matrices, which refer to the marginals that are chosen to become noisy measurements as well as the PLB allocation for each marginal, which in turn defines the scale of the noise added to the confidential CEF-based marginal, for each geographic level. As described in [abowd20222020] in more detail, we use the $\rho$-zero-concentrated differential privacy ($\rho$-zCDP) accounting framework to specify these PLB values [bun2016concentrated]. Each strategy matrix described in the next two subsections provides the initial PLB allocations for each geographic unit on the spine, prior to the spine optimization routines’ alteration of these allocations for parent and child geographic units for which the parent has only one child; for more detail on the spine optimization routines, see[cumings2022geographic]. Aside from dictating the scale of the noise for each noisy measurement used within the DAS, the PLB allocations are also useful because they can be used to derive upper bounds on the power of certain hypothesis tests on any given individual record in the CEF; for more detail, see [kifer2022bayesian].